OpenStudy (anonymous):

Help!! Can you solve this and explain the properties used? log9x+log9(x-8)=1

OpenStudy (anonymous):

log(x/y) = log(x) - log(y) log(x*y) = log(x) + log(y) log(x^2) = 2log(x)

OpenStudy (anonymous):

first step is to distribute the 9 in log9(x-8)

OpenStudy (anonymous):

What does that result in?

OpenStudy (anonymous):

Log9x (X^2-8x)

OpenStudy (anonymous):

log9x+log9(x-8) = 1 ==> log((9x)(9x-72)) = 1

Nnesha (nnesha):

9 is base \[\large\rm log_9~ x + \log_9 ~(x-8) =1\]

OpenStudy (anonymous):

oooh!

Nnesha (nnesha):

use the rules which is \[\huge\rm log x + \log y = \log (x \times y )\] like manumben gve u

Nnesha (nnesha):

this*

OpenStudy (anonymous):

Right, and then what do I do?

OpenStudy (anonymous):

log_a(X) = b means a^b = X

OpenStudy (anonymous):

where a is the base

Nnesha (nnesha):

\[\huge\rm log_9 ~ x + \log_9 (x-8) \] there is plus sign so you have to change this to multiplication remember like this log x + log y = log ( x times y ) theere is common base so log_9 ??????

OpenStudy (anonymous):

Right and I used the Product property and got Log9 (x^2-8)

OpenStudy (anonymous):

And so my equation now is Log9(x^2-8)=1 and I don't know where to go from there

Nnesha (nnesha):

x(x-8 ) = ???

OpenStudy (anonymous):

I multiplied that through which gave me x^2-8x

Nnesha (nnesha):

yes 8x is right now change log to exponential form